The National Transportation Safety Bureau indicates that a person in a car crash has a reasonable chance of survival if his or her acceleration is less than 300 m/s2. Suppose that the mass of a car is 1200 kg .

See rest of question below in comments:
Answer #1

a) the deceleration required is 300 m/s^2 (as the person inside the car will be accelerated by same amount of force due to inertia)

There fore the for requiredvto stop F= ma

= 1200*300

=360000

=360kN

b) v^2 - u^2 = 2* a* s

u= 21 m/s v=0

-21^2 = -2*300*S

S= 42/600

S=0.07 m